# Generating Molecular Conformer Fields

### v2

@@ Abstract
In this paper we tackle the problem of generating conformers of a molecule in 3D space given its molecular graph. We parameterize these conformers as continuous functions that map elements from the molecular graph to points in 3D space. We then formulate the problem of learning to generate conformers as learning a distribution over these functions using a diffusion generative model, called Molecular Conformer Fields (MCF). Our approach is simple and scalable, and achieves state-of-the-art performance on challenging molecular conformer generation benchmarks while making no assumptions about the explicit structure of molecules (e.g. modeling torsional angles). MCF represents an advance in extending diffusion models to handle complex scientific problems in a conceptually simple, scalable and effective manner.
@@ 1 Introduction
In this paper we tackle the problem of Molecular Conformer Generation, i.e.  predicting the diverse low-energy three-dimensional conformers of molecules, relying solely on their molecular graphs as illustrated in Fig. 1. Molecular Conformer Generation is a fundamental problem in computational drug discovery and chemo-informatics, where understanding the intricate interactions between molecular and protein structures in 3D space is critical, affecting aspects such as charge distribution, potential energy, etc. (Batzner et al., 2022). The core challenge associated with conformer generation springs from the vast complexity of the 3D structure space, encompassing factors such as bond lengths and torsional angles. Despite the molecular graph dictating potential 3D conformers through specific constraints, such as bond types and spatial arrangements determined by chiral centers, the conformational space experiences exponential growth with the expansion of the graph size and the number of rotatable bonds (Axelrod & Gomez-Bombarelli, 2022). This complicates brute force approaches, making them virtually unfeasible for even moderately small molecules.
Figure 1: We formulate the molecular conformer generation problem as learning a distribution of functions over graphs. Each conformer is a field (e.g. a function) $f$ that can map elements from a particular molecular graph $\mathcal{G}$ to points in 3D space, $f:\mathcal{G}\rightarrow\mathbb{R}^{3}$. We call these conformer fields as they can be evaluated for any point in $\mathcal{G}$. We visualize an example of example of this conformer field where points are colored based on the distance to their closest atom (see Sect. A.5 for details).
@@
Systematic methods, like OMEGA (Hawkins et al., 2010), offer rapid processing through rule-based generators and curated torsion templates. Despite their efficiency, these models typically fail on complex molecules, as they often overlook global interactions and are tricky to extend to inputs like transition states or open-shell molecules. Classic stochastic methods, like molecular dynamics (MD) and Markov chain Monte Carlo (MCMC), rely on extensively exploring the energy landscape to find low-energy conformers. Such techniques suffer from sampling inefficiency for large molecules and struggle to generate diverse representative conformers (Hawkins, 2017; Wilson et al., 1991; Grebner et al., 2011). In the domain of learning-based approaches, several works have looked at conformer generation problems through the lens of probabilistic modeling, using either normalizing flows (Xu et al., 2021a) or diffusion models (Xu et al., 2022; Jing et al., 2022). These approaches tend to use equivariant network architectures to deal with molecular graphs (Xu et al., 2022) or model domain-specific factors like torsional angles (Ganea et al., 2021; Jing et al., 2022). However, explicitly enforcing these domain-specific inductive biases can sometimes come at a cost.For example, Torsional Diffusion relies on rule-based methods to find rotatable bonds which may fail especially for complex molecules. Also, the quality of generated conformers are adhered to the non-differentiable cheminformatic methods used to predict local substructures. On the other hand, recent works have proposed domain-agnostic approaches for generative modeling of data in function space (Du et al., 2021; Dupont et al., 2022b, a; Zhuang et al., 2023) obtaining great performance. As an example, in (Zhuang et al., 2023) the authors use a diffusion model to learn a distribution over fields $f$, showing great results on different data domains like images (i.e.  $\ f:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$) or 3D geometry (i.e.  $\ f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{1}$), where the domain of the function $\mathbb{R}^{n}$ is fixed across functions. However, dealing with fields defined on different domains (e.g. different molecular graphs, as in molecular conformer generation) still remains an open problem.
To address these issues, we present Molecular Conformer Fields (MCF), an approach to learn generative models of molecular conformers. We interpret conformers as fields/functions (we use both terms exchangeably) on graphs that map elements in the graph $\mathcal{G}_{i}$ to points in $\mathbb{R}^{3}$, $f_{i}:\mathcal{G}_{i}\rightarrow\mathbb{R}^{3}$, which we define as conformer fields. MCF present a generalization over previous work on generative models in function space (Du et al., 2021; Dupont et al., 2022b, a; Zhuang et al., 2023) to deal with functions $f_{i}$ defined on different graphs $\mathcal{G}_{i}$ while also generalizing those approaches to dealing with intrinsic coordinate systems. MCF outperforms recent conformer generation benchmarks while being conceptually simpler to understand and efficient to scale. In addition, successfully tackling the problem of conformer generation with a general approach like MCF shows that it is a strong method that can generalize to other scientific disciplines with little effort. In addition, MCF can be a stepping stone towards all-purpose diffusion generative models.
Our contributions can be summarized as follows:
  • We provide a simple yet effective approach for molecule conformer generation that achieves state-of-the-art performance on standard benchmarks.
@@
  • Our approach directly predicts the 3D position of atoms as opposed to torsional angles or other domain-specific variables, providing a simple and scalable training recipe.
  • We provide an extensive ablation study to understand what are the critical factors that affect the performance of our proposed MCF in molecular conformer generation.
@@ 2 Related Work
Recent works have tackled the problem of conformer generation using learning-based generative models. (Simm & Hernández-Lobato, 2019) and (Xu et al., 2021b) develop two-stage methods which first generate interatomic distances following VAE framework and then predict conformers based on the distances. (Guan et al., 2021) propose neural energy minimization to optimize low-quality conformers. In (Xu et al., 2021a), a normalizing flow approach is proposed as an alternative to VAEs. To avoid the accumulative errors from two-stage generation, (Shi et al., 2021) implement score-based generative model to directly model the gradient of logarithm density of atomic coordinates. In GeoDiff (Xu et al., 2022), a diffusion model is used which focuses on crafting equivariant forward and backward processes with equivariant graph neural networks. In GeoMol (Ganea et al., 2021), the authors first predict 1-hop local structures and then propose a regression objective coupled with an Optimal Transport loss to predict the torsional angles that assemble substructures of a molecule. Following this, Torsional Diffusion (Jing et al., 2022) proposed a diffusion model on the torsional angles of the bonds rather than a regression model used in (Ganea et al., 2021).
Our approach extends recent efforts in generative models for functions in Euclidean space (Zhuang et al., 2023; Dupont et al., 2022b, a; Du et al., 2021), to functions defined over graphs (e.g. chemical structure of molecules). The term Implicit Neural Representation (INR) is used in these works to denote a parameterization of a single function (e.g.  a single image in 2D) using a neural network that maps the function’s inputs (i.e.   pixel coordinates) to its outputs (i.e.  RGB values). Different approaches have been proposed to learn distributions over fields in Euclidean space; GASP (Dupont et al., 2022b) leverages a GAN whose generator produces field data whereas a point cloud discriminator operates on discretized data and aims to differentiate real and generated functions. Two-stage approaches (Dupont et al., 2022a; Du et al., 2021) adopt a latent field parameterization (Park et al., 2019) where functions are parameterized via a hyper-network (Ha et al., 2017) and a generative model is learnt on the latent or INR representations. In addition, our approach also relates to recent work focusing on fitting a function (e.g. learning an INR) on a manifold using an intrinsic coordinate system (Koestler et al., 2022; Grattarola & Vandergheynst, 2022), and generalizes it to the problem of learning a probabilistic model over multiple functions defined on different manifolds/graphs. Intrinsic coordinate systems have also been used in Graph Transformers to tackle supervised learning tasks (Maskey et al., 2022; Sharp et al., 2022; He et al., 2022; Dwivedi et al., 2020).
@@ 3.1 Diffusion Probabilistic Fields
Diffusion Probabilistic Fields (DPF) (Zhuang et al., 2023) belong to the broad family of latent variable models (Everett, 2013) and can be consider a generalization of DDPMs (Ho et al., 2020) to deal with functions $f:\mathbf{M}\rightarrow Y$ which are infinite dimensional. Conceptually speaking, DPF (Zhuang et al., 2023) parameterizes functions $f$ with a set of context pairs containing input-outputs to the function. Using these context pairs as input to DPF, the model is trained to denoise any continuous query coordinate (e.g. query pairs) in the domain of the function at timestep $t$. In order to learn a parametric distribution over functions $p_{\theta}(f_{0})$ from an empirical distribution of functions s $q(f_{0})$, DPF reverses a diffusion Markov Chain that generates function latents $f_{1:T}$ by gradually adding Gaussian noise to (context) input-output pairs randomly drawn from $f\sim q(f_{0})$ for $T$ time-steps as follows: $q(f_{t}|f_{t-1}):=\mathcal{N}\left(f_{t-1};\sqrt{\bar{\alpha}_{t}}f_{0},(1-\bar{\alpha}_{t}){\mathbf{I}}\right)$. Here, $\bar{\alpha}_{t}$ is the cumulative product of fixed variances with a handcrafted scheduling up to time-step $t$. DPF (Zhuang et al., 2023) follows the training recipe in (Ho et al., 2020) in which: i) The forward process adopts sampling in closed form. ii) reversing the diffusion process is equivalent to learning a sequence of denoising (or score) networks $\epsilon_{\theta}$, with tied weights. Reparameterizing the forward process as $f_{t}=\sqrt{\bar{\alpha}_{t}}f_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ results in the “simple” DDPM loss: $\mathbb{E}_{t\sim[0,T],f_{0}\sim q(f_{0}),\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}f_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}\right]$, which makes learning of the data distribution $p_{\theta}(f_{0})$ both efficient and scalable. At inference time, DPF computes $f_{0}\sim p_{\theta}(f_{0})$ via ancestral sampling (Zhuang et al., 2023). Concretely, DPF starts by sampling dense query coordinates and assigning a gaussian value to them $f_{T}\sim\mathcal{N}({\bm{0}},\mathbf{I})$. Then, it iteratively applies the score network $\epsilon_{\theta}$ to denoise $f_{T}$, thus reversing the diffusion Markov Chain to obtain $f_{0}$. In practice, DPFs have obtained amazing results for signals living in an Euclidean geometry. However, the extension to functions defined on graphs remains an open problem.
@@ 3.2 Conformers as Functions on Graphs
Following the setting in previous work (Xu et al., 2022; Ganea et al., 2021; Jing et al., 2022) a molecule with $n$ atoms is represented as an undirected graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$, where $\mathcal{V}=\{v_{i}\}^{n}_{i=1}$ is the set of vertices representing atoms and =$\mathcal{E}=\{e_{ij}|(i,j)\subseteq|\mathcal{V}|\times|\mathcal{V}|\}$ is the set of edges representing inter-atomic bonds. In this paper, we parameterize a molecule’s conformer as a function $f:\mathcal{G}\rightarrow\mathbb{R}^{3}$ that takes atoms in the molecular graph $\mathcal{G}$ and maps them to 3D space, we call this function a conformer field. The training set is composed of conformer fields $f_{i}:\mathcal{G}_{i}\rightarrow\mathbb{R}^{3}$, that each maps atoms of a different molecule $\mathcal{G}_{i}$ to a 3D point. We then formulate the task of conformer generation as learning a prior over a training set of conformer fields. We drop the subscript $i$ in the remainder of the text for notation simplicity.
@@ 3.3 Equivariance in Conformer Generation
Equivariance has become an important topic of study in generative models (Abbott et al., 2023, 2022; Kanwar et al., 2020). In particular, enforcing equivariance as an explicit inductive bias in neural networks can often lead to improved generalization (Köhler et al., 2020) by constraining the space of functions that can be represented by a model. On the other hand, recent literature shows that models that can learn these symmetries from data rather than explicitly enforcing them (e.g. transformers vs CNNs) tend to perform better as they are more amenable to optimization (Bai et al., 2021).
Equivariance also plays an interesting role in conformer generation. On one hand, it is important when training likelihood models of conformers, as the likelihood of a conformer is invariant to roto-translations (Köhler et al., 2020). On the other hand, when training models to generate conformers given a molecular graph, explicitly baking roto-translation equivariance might not be as necessary. This is because the intrinsic structure of the conformer encodes far more information about its properties than the extrinsic coordinate system (eg. rotation and translation) in which the conformer is generated (Ruddigkeit et al., 2012). In addition, recent approaches for learning simulations on graphs (Sanchez-Gonzalez et al., 2020) or pre-training models for molecular prediction tasks (Zaidi et al., 2022) have relied on non-equivariant architectures.
We follow this trend and empirically show that explicitly enforcing roto-translation equivariance is not a strong requirement for generalization. Furthermore, we also show that approaches that do not explicitly enforce roto-translation equivariance (like ours) can match or outperform approaches that do.

### v3

# Swallowing the Bitter Pill: Simplified Scalable Conformer Generation

@@ Abstract
We present a novel way to predict molecular conformers through a simple formulation that sidesteps many of the heuristics of prior works and achieves state of the art results by using the advantages of scale. By training a diffusion generative model directly on 3D atomic positions without making assumptions about the explicit structure of molecules (e.g. modeling torsional angles) we are able to radically simplify structure learning, and make it trivial to scale up the model sizes. This model, called Molecular Conformer Fields (MCF), works by parameterizing conformer structures as functions that map elements from a molecular graph directly to their 3D location in space. This formulation allows us to boil down the essence of structure prediction to learning a distribution over functions. Experimental results show that scaling up the model capacity leads to large gains in generalization performance without enforcing inductive biases like rotational equivariance. MCF represents an advance in extending diffusion models to handle complex scientific problems in a conceptually simple, scalable and effective manner.
@@ 1 Introduction
In this paper we tackle the problem of molecular conformer generation, i.e.  predicting the diverse low-energy three-dimensional conformers of molecules. Molecular conformer generation is a fundamental problem in computational drug discovery and chemo-informatics, where understanding the intricate interactions between molecular and protein structures in 3D space is critical, affecting aspects such as charge distribution, potential energy, etc. (Batzner et al., 2022). The core challenge associated with conformer generation is the vast complexity of the 3D structure space, encompassing factors such as bond lengths and torsional angles. Despite the molecular graph dictating potential 3D conformers through specific constraints, such as bond types and spatial arrangements determined by chiral centers, the conformational space experiences exponential growth with the expansion of the graph size and the number of rotatable bonds (Axelrod & Gomez-Bombarelli, 2022). This complicates brute force and exhaustive approaches, making them virtually unfeasible for even moderately small molecules.
Figure 1: Overview of proposed MCF. The structure of molecular graph is encoded through eigenvectors of Laplacian eigen-decomposition $\varphi(\mathcal{V})$ and atomic features $\mathcal{A}$. MCF directly operates on atom coordinates in 3D space and trains the diffusion model to denoise the function in 3D coordinates. The score network is developed with attention-based PerceiverIO architecture. Context pairs $\mathbf{C}_{t}$ attend to a latent array of learnable parameters via cross attention and the latent array goes through several self attention blocks. Finally, the query pairs $\mathbf{Q}_{t}$ cross-attend to the latent array to produce the final noise prediction $\hat{\epsilon}_{q}$ in 3D space.
@@
Systematic methods, like OMEGA (Hawkins et al., 2010), offer rapid processing through rule-based generators and curated torsion templates. Despite their efficiency, these models typically fail on complex molecules, as they often overlook global interactions and are tricky to extend to inputs like transition states or open-shell molecules. Classic stochastic methods, like molecular dynamics (MD) and Markov chain Monte Carlo (MCMC), rely on extensively exploring the energy landscape to find low-energy conformers. Such techniques suffer from sampling inefficiency for large molecules and struggle to generate diverse representative conformers (Hawkins, 2017; Wilson et al., 1991; Grebner et al., 2011). In the domain of learning-based approaches, several works have looked at conformer generation problems through the lens of probabilistic modeling, using either normalizing flows (Xu et al., 2021a) or diffusion models (Xu et al., 2022; Jing et al., 2022). These approaches tend to use equivariant network architectures to deal with molecular graphs (Xu et al., 2022) or model domain-specific factors like torsional angles (Ganea et al., 2021; Jing et al., 2022). However, explicitly enforcing these domain-specific inductive biases come at a cost. For example, Torsional Diffusion models rely on rule-based methods to find rotatable bonds which may fail especially for complex molecules. Ultimately, the quality of generated conformers is destined to suffer from errors of the non-differentiable cheminformatic methods used to predict local substructures. On the other hand, recent works have proposed domain-agnostic approaches for generative modeling of data in function space (Du et al., 2021; Dupont et al., 2022b, a; Zhuang et al., 2023) obtaining great performance. As an example, Zhuang et al. (2023) use a diffusion model to learn a distribution over functions $f$, showing great results on different data domains like images (i.e.  $\ f:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$) or 3D geometry (i.e.  $\ f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{1}$), where the domain of the function $\mathbb{R}^{n}$ is fixed across functions. Such frameworks provide a valuable paradigm to investigate whether domain-agnostic methods with little to no inductive biases can be successfully transferred to solve scientific problems (e.g. molecular conformer generation).
To this end, we present Molecular Conformer Fields (MCF), a simple and scalable approach to learn generative models of molecular conformers. We leverage a domain-agnostic architecture that makes no assumptions about molecular structures and trivially benefits from scale. We formulate the molecular conformer generation problem as learning a distribution over functions/fields (we use both terms exchangeably), an approach that has been applied widely to various data domains (Zhuang et al., 2023). Specifically, conformers are interpreted as functions that map points on graph $\mathcal{G}_{i}$ to atom coordinates in $\mathbb{R}^{3}$, $f_{i}:\mathcal{G}_{i}\rightarrow\mathbb{R}^{3}$, which we call a conformer field. Unlike many prior efforts that shoe-horn inductive biases of molecular structures into the model (e.g. developing equivariant diffusion process, modeling torsional angles, etc.) (Xu et al., 2022; Ganea et al., 2021; Jing et al., 2022), MCF operates directly on 3D atom coordinates, without enforcing molecular constraints explicitly, letting the model learn these directly from the data.
Instead of using Graph Neural Networks with intricate equivariance designs, MCF builds a score network using PerceiverIO (Jaegle et al., 2022) (see Fig. 1) which is a scalable and efficient variant of the Transformer architecture. Our model is simple to implement and efficient to scale. Experiments on recent conformer generation benchmarks show MCF surpasses strong baselines by a gap that gets larger as we scale model capacity, potentially revealing a bitter lesson (Sutton, 2019) moment for conformer generation, when large models with fewer domain-specific architectural inductive biases lead to better performance. Superior performance of MCF on molecular conformation generation highlights the potential for building a singe domain-agnostic method that is simple and scalable to work on many different problems.
Our contributions are summarized as follows:
  • We introduce a novel approach for molecular conformer generation that has strong scaling properties and surpasses previous methods by a large margin on standard benchmarks.
@@
  • Our approach directly predicts the 3D position of atoms as opposed to domain-specific variables, providing a simple and scalable training recipe.
  • MCF shows that enforcing inductive biases like rotational equivariance or modeling torsional angles is not required for generalization.
@@ 2 Related Work
Recent works have tackled the problem of molecular conformer generation using learning-based generative models. Simm & Hernández-Lobato (2019) and Xu et al. (2021b) develop two-stage methods which first generate interatomic distances following VAE framework and then predict conformers based on the distances. Guan et al. (2021) propose neural energy minimization to optimize low-quality conformers. In Xu et al. (2021a), a normalizing flow approach is proposed as an alternative to VAEs. To avoid the accumulative errors from two-stage generation, Shi et al. (2021) implement score-based generative model to directly model the gradient of logarithm density of atomic coordinates. In GeoDiff (Xu et al., 2022), a diffusion model is used which focuses on crafting equivariant forward and backward processes with equivariant graph neural networks. In GeoMol (Ganea et al., 2021), the authors first predict 1-hop local structures and then propose a regression objective coupled with an Optimal Transport loss to predict the torsional angles that assemble substructures of a molecule. Following this, Torsional Diffusion (Jing et al., 2022) proposed a diffusion model on the torsional angles of the bonds rather than a regression model used in Ganea et al. (2021).
Our approach extends recent efforts in generative models for functions in Euclidean space (Zhuang et al., 2023; Dupont et al., 2022b, a; Du et al., 2021), to functions defined over graphs (e.g. chemical structure of molecules). Different approaches have been proposed to learn distributions over fields in Euclidean space; GASP (Dupont et al., 2022b) leverages a GAN whose generator produces field data whereas a point cloud discriminator operates on discretized data and aims to differentiate real and generated functions. Two-stage approaches (Dupont et al., 2022a; Du et al., 2021) adopt a latent field parameterization (Park et al., 2019) where functions are parameterized via a hyper-network (Ha et al., 2017) and a generative model is learnt in latent space. MCF presents a generalization over these approaches to deal with training sets where each function $f_{i}$ is defined on a different graph $\mathcal{G}_{i}$, as opposed to in Euclidean space. In addition, MCF also related to recent work focusing on fitting a function on a manifold using an intrinsic coordinate system (Koestler et al., 2022; Grattarola & Vandergheynst, 2022), and generalizes it to the problem of learning a probabilistic model over multiple functions defined on different graphs. Intrinsic coordinate systems have also been used in Graph Transformers to tackle supervised learning tasks (Maskey et al., 2022; Sharp et al., 2022; He et al., 2022; Dwivedi et al., 2020).
@@ 3.1 Diffusion Probabilistic Fields
Diffusion Probabilistic Fields (DPF) (Zhuang et al., 2023) belongs to the broad family of latent variable models (Everett, 2013) and can be consider a generalization of DDPMs (Ho et al., 2020) to deal with functions $f:\mathbf{M}\rightarrow Y$ which are infinite dimensional. Conceptually speaking, DPF (Zhuang et al., 2023) parameterizes functions $f$ with a set of context pairs containing input-outputs to the function. Using these context pairs as input to DPF, the model is trained to denoise any query coordinate (e.g. query pairs) in the domain of the function at timestep $t$ (as shown in Fig. 1). In order to learn a parametric distribution over functions $p_{\theta}(f_{0})$ from an empirical distribution of functions s $q(f_{0})$, DPF reverses a diffusion Markov Chain that generates function latents $f_{1:T}$ by gradually adding Gaussian noise to (context) input-output pairs randomly drawn from $f\sim q(f_{0})$ for $T$ time-steps as follows: $q(f_{t}|f_{t-1}):=\mathcal{N}\left(f_{t-1};\sqrt{\bar{\alpha}_{t}}f_{0},(1-\bar{\alpha}_{t}){\mathbf{I}}\right)$. Here, $\bar{\alpha}_{t}$ is the cumulative product of fixed variances $\alpha_{t}$ with a handcrafted scheduling up to time-step $t$. DPF (Zhuang et al., 2023) follows the training recipe in Ho et al. (2020) in which: i) The forward process adopts sampling in closed form. ii) reversing the diffusion process is equivalent to learning a sequence of denoising (or score) networks $\epsilon_{\theta}$, with tied weights. Reparameterizing the forward process as $f_{t}=\sqrt{\bar{\alpha}_{t}}f_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ results in the “simple” DDPM loss: $\mathbb{E}_{t\sim[0,T],f_{0}\sim q(f_{0}),\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}f_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}\right]$, which makes learning of the data distribution $p_{\theta}(f_{0})$ both efficient and scalable. At inference time, DPF computes $f_{0}\sim p_{\theta}(f_{0})$ via ancestral sampling (Zhuang et al., 2023). Concretely, DPF starts by sampling dense query coordinates and assigning a gaussian value to them $f_{T}\sim\mathcal{N}({\bm{0}},\mathbf{I})$. Then, it iteratively applies the score network $\epsilon_{\theta}$ to denoise $f_{T}$, thus reversing the diffusion Markov Chain to obtain $f_{0}$. In practice, DPFs have obtained amazing results for signals living in an Euclidean geometry.
@@ 3.2 Conformers as Functions on Graphs
Following the setting in previous work (Xu et al., 2022; Ganea et al., 2021; Jing et al., 2022) a molecule with $n$ atoms is represented as an undirected graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$, where $\mathcal{V}=\{v_{i}\}^{n}_{i=1}$ is the set of vertices representing atoms and =$\mathcal{E}=\{e_{ij}|(i,j)\subseteq|\mathcal{V}|\times|\mathcal{V}|\}$ is the set of edges representing inter-atomic bonds. We further use $\mathcal{A}$ to denote atomic features which also are leveraged by our generative model. In this paper, we parameterize a molecule’s conformer as a function $f:\mathcal{G}\rightarrow\mathbb{R}^{3}$ that takes atoms (e.g. vertices) in the molecular graph $\mathcal{G}$ and maps them to 3D space, we call this function a conformer field. The training set is composed of conformer fields $f_{i}:\mathcal{G}_{i}\rightarrow\mathbb{R}^{3}$, where each field maps atoms of a different molecule $\mathcal{G}_{i}$ to a 3D point. We then formulate the task of conformer generation as learning a prior over a training set of conformer fields. We drop the subscript $i$ in the remainder of the text for notation simplicity.
@@ 3.3 Equivariance in Conformer Generation
Equivariance has become an important topic of study in generative models (Abbott et al., 2023, 2022; Kanwar et al., 2020). In particular, enforcing equivariance as an explicit inductive bias in neural networks can lead to improved generalization (Köhler et al., 2020) by constraining the space of functions that can be represented by a model. On the other hand, recent literature shows that models that can learn these symmetries from data rather than explicitly enforcing them (e.g. Transformers vs CNNs) tend to perform better as they are more amenable to optimization (Bai et al., 2021).
Equivariance also plays an interesting role in conformer generation. On one hand, it is important when training likelihood models of conformers, as the likelihood of a conformer is invariant to roto-translations (Köhler et al., 2020). On the other hand, when training models to generate conformers given a molecular graph, explicitly baking roto-translation equivariance might not be as necessary. This is because the intrinsic structure of the conformer encodes far more information about its properties than the extrinsic coordinate system (eg. rotation and translation) in which the conformer is generated (Ruddigkeit et al., 2012). In addition, recent approaches for learning simulations on graphs (Sanchez-Gonzalez et al., 2020) or pre-training models for molecular prediction tasks (Zaidi et al., 2022) have successfully relied on non-equivariant architectures.
In this paper, we ask whether inductive biases like rotational equivariance can be traded for model scale in general purposes architectures like Transformers. Our empirical results show that explicitly enforcing roto-translation equivariance is not a strong requirement for generalization. Furthermore, we show that scalable approaches that do not explicitly enforce roto-translation equivariance (like ours) can outperform approaches that do by a large margin .